OpenStudy (sparklestaraa):

Choose the equation below that represents the line passing through the point (-5, 1) with a slope of 3/2

OpenStudy (sparklestaraa):

@misty1212

OpenStudy (misty1212):

HI!

OpenStudy (misty1212):

they all got \(\frac{3}{2}\)

OpenStudy (misty1212):

use the point slope formula \[y-y_1=m(x-x_1)\] and you get \[y-1=\frac{3}{2}(x-(-5))\] or just \[y-1=\frac{3}{2}(x+5)\]

OpenStudy (sparklestaraa):

: D